### v1

# Vision-and-Language Training Helps Deploy Taxonomic Knowledge but Does Not Fundamentally Alter It

Abstract
Does vision-and-language (VL) training change the linguistic representations of language models in meaningful ways? Most results in the literature have shown inconsistent or marginal differences, both behaviorally and representationally. In this work, we start from the hypothesis that the domain in which VL training could have a significant effect is lexical-conceptual knowledge, in particular its taxonomic organization. Through comparing minimal pairs of text-only LMs and their VL-trained counterparts, we first show that the VL models often outperform their text-only counterparts on a text-only question-answering task that requires taxonomic understanding of concepts mentioned in the questions. Using an array of targeted behavioral and representational analyses, we show that the LMs and VLMs do not differ significantly in terms of their taxonomic knowledge itself, but they differ in how they represent questions that contain concepts in a taxonomic relation vs. a non-taxonomic relation. This implies that the taxonomic knowledge itself does not change substantially through additional VL training, but VL training does improve the deployment of this knowledge in the context of a specific task, even when the presentation of the task is purely linguistic.
1 Introduction
Humans readily integrate perceptual and linguistic signals to form generalizable mappings from semantic information to language, allowing them to reason about concepts beyond their immediate environment [56, 17]. Approaches to concept grounding in AI, which traditionally relied on annotated datasets to specify how language links to people, objects, and events [70, 25], have rapidly shifted in light of the impressive capabilities of vision-language models (VLMs). footnotetext: ${}^{*},^{\dagger}$: Equal contribution; Code can be found at https://github.com/tinlaboratory/taxonomigqa
Many standard VLMs [32, 29, i.a.] often build on top of a pretrained language model (LM) by adding visual conditioning to its next token prediction task, often also updating the parameters of the language model. Analyses of VLM capabilities often focus on the multimodal tasks this additional modality enables. But (how) does this vision-and-language (VL) training change the linguistic capacity of the model? Answering this question requires comparing VL-tuned LMs to their original LM counterparts. Empirical evidence in this literature is rather sparse, often comparing such “VLM-LM minimal pairs” on general benchmarks such as MMLU [16] and GLUE [63]. In this paper, we consider a more targeted investigation (like [71]) of VLM-LM pairs in a particular domain: lexical-conceptual knowledge, specifically its taxonomic organization (e.g., a cat is an animal). Evaluation of taxonomic knowledge has been of continued interest within the Natural Language Processing [14, 31, 42, 44] and Computer Vision communities [2, 61, 47]—however, to the best of our knowledge no work so far has compared minimally differing VLM-LM pairs in terms of how well they can reason taxonomically.
To this end, we develop TaxonomiGQA, a synthetically augmented text-only version of the popular visual-question answering (VQA) dataset GQA [18], where a subset of WordNet [39] hierarchy is used to create questions that require taxonomic knowledge. On comparing 7 widely used VLM-LM minimal pairs, we find most VLMs to consistently outperform their LM counterparts, despite the fact that the QA task is text only. We put forth two hypotheses to explain these results. H1: VL training fundamentally alters the (task-agnostic) taxonomic knowledge in LMs; and H2: VL training improves the ability of the LM to deploy its (largely unchanged) taxonomic knowledge in tasks that require its usage. Through a series of controlled behavioral and representational analyses, we find evidence that supports H2 relative to H1. Finally, we conduct a preliminary investigation where we relate the successes of VLMs over LMs to the visual similarities between the hyponym-hypernym categories we have tested in our work. Here we find initial evidence that suggests that VLMs especially perform well at answering questions about hyponym-hypernym pairs that are visually similar, leaving open areas of interesting future research for a more precise characterization of the role of visual input.
2 Related Work
Influence of vision on language in VLMs
There are two main strands of empirical work measuring the influence of the additional visual modality on models’ linguistic behavior and representations. The first line of work compares VLM and LM performance on downstream text-only benchmarks. The results are mixed: for instance, FLAVA [59] noted around 8% point gains over the base masked language model on GLUE-style NLP tasks (although the evaluation setting involved finetuning). On the other hand, Molmo has been reported to be outperformed by its base LM, Qwen, on text-only benchmarks like MMLU [8]. Generally, more evidence exists in favor of multimodal training hurting text-only task performance [20, 36] and this observation has been used to argue for freezing the language part of the model during multimodal training [12]. The second line of work conducts more targeted comparisons of VLMs and LMs, examining whether additional vision training leads to differences in representations of syntactic categories [64] and performance on tasks that require more “grounding” [71], but the findings overall have indicated no substantial differences.
Taxonomic knowledge and its deployment
Taxonomic knowledge has long been a topic of interest in cognitive psychology [34, 45], and has also often been used to analyze conceptual organization in LMs [14, 31, 44, 43]. Work that tests its functional consequences, such as property inheritance [42, 57, 55] and inductive generalization [41, 13], found strong evidence that while LMs do learn explicit taxonomic knowledge, they struggle to deploy it in taxonomically sensitive tasks [42]. Taxonomic knowledge has also been evaluated and analyzed in multimodal models. For instance, Pach et al. [47] show that the internal structure of neurons in models such as CLIP are often in alignment with existing taxonomies. Our work contributes to this line of work by proposing a level-ground comparison between minimally differing LMs and VLMs, narrowing in on the precise ways in which additional VL training may or may not alter the nature of this knowledge.
Figure 1: The three-step pipeline to create TaxonomiGQA.
3 Behavioral testing of minimal pair VLMs and LMs with TaxonomiGQA
The question we are interested in answering concerns the change that VL training introduces to the lexical-conceptual knowledge of a model. This requires a shared evaluation that can be applied to both VLMs and LMs. We discuss below how we designed this evaluation as well as our findings about a range of VLM-LM pairs from this evaluation.
3.1 Dataset design
We created a QA dataset that requires taxonomic understanding based on GQA [18], named TaxonomiGQA. A datapoint in GQA consists of an image of a scene, a question about this scene, and metadata that includes a scene graph of the objects, their attributes, and relations between the objects. We applied a three-step modification (each step illustrated in Figure 1) to create a text-only dataset, since our goal is to systematically compare the LM and VLMs’ taxonomic competence. (1) Convert the scene graph into a purely textual description of the scene programmatically using hand-crafted templates; (2) For each question that contains a word that corresponds to a node in our reference taxonomy, substitute the word to its hypernym; (3) For each substitution, create four negative samples following Misra et al. [42], by substituting the target word with a word that is not in its hypernymy chain. Below we describe the resources, filtering, and sampling details used to create TaxonomiGQA.
Taxonomy
To construct the reference taxonomy, we first extracted all unique noun lemmas ($N=1216$) that appeared in the GQA questions, and annotated their senses in the WordNet taxonomy, these serve as the leaf nodes of our taxonomy. Next, for each noun, we extracted its hypernym chain (e.g., dog < canine < mammal < vertebrate < animal) from WordNet, rejecting hypernyms that were too abstract (determined manually), e.g., entity, material, conveyance. 315 concepts were removed as a result, many of which often had abstract entities in their hypernym chains or had non-ideal WordNet categorization (e.g., bubble as a member of ball), leaving us with 901 unique chains.
Dataset construction
We applied a multi-stage filtering process to the validation split of GQA (10,696 images/scenes and 488,293 questions) to obtain our base questions. We first applied scene-level filtering by excluding scenes containing more than 20 annotated objects or any repeated object labels to avoid ambiguity in referring expressions in text. For each remaining scene, we applied question-level filtering to retain questions that refer to a single object (excluding any that mention multiple objects) and whose hypernyms do not overlap with those of any other object in the scene. Next, we balanced the dataset by randomly sampling 40 questions per scene in proportion to each scene’s question type ratios. We further filtered the questions by answer type and restricted the dataset to yes/no questions to facilitate the substitution step. This reduced our taxonomy to 314 unique chains. In the base questions remaining after filtering, we substituted each target concept with each of its hypernyms in its hypernym chain to obtain the substituted questions. Then, we created negative sample questions by substituting the target concepts with concepts that are not in their hypernym chain, discarding question types where this substitution was not possible due to the introduction of presupposition failure (e.g., questions such as Is the color of the dog brown? when there is no dog in the image). This ended up eliminating more hypernym chains (which were only present in the discarded question types), leaving us with 126 final chains. More details about this negative sampling pipeline is given in Appendix C.
Dataset statistics
The final dataset contains 1,342 unique images/scenes, 29,604 positive sample instances (9,334 targeting leaf node concepts, 20,270 targeting hypernym-substitutions), and 4 negative samples for each positive sample, amounting to 148,020 total instances. There are 276 hyponym-hypernym pairs, 126 unique hypernym chains, 88 unique hypernyms, and 24 top-level categories.
3.2 Metrics
We propose metrics designed to be sensitive to taxonomic structure (cf. [61]). The design principles are: (1) be sensitive to hierarchical relationships between two concepts; (2) anchor expectations on taxonomic knowledge conditioned on the model’s success at foundational or prerequisite tasks; and (3) provide insight into robustness, including contrasting the performance on both positive and negative samples. By grounding our metrics in these properties, we move beyond correctness and toward a more systematic assessment of model performance.
[…]
Overall Accuracy
measures the proportion of time the model correctly answers all original, unsubstituted, and hypernym-substituted instances, treating each instance as separate item.
$\text{Overall}=\frac{1}{N+\sum_{i=1}^{N}k_{i}}\left[\sum_{i=1}^{N}\left(\texttt{correct}(X_{i})+\sum_{j=1}^{k_{i}}\texttt{correct}(X^{s}_{j})\right)\right]$ (1)
Conditional Accuracy
measures the proportion of time the model correctly answers hypernym-substituted instances, conditioned on the fact that the model correctly answered the original, unsubstituted instance correctly. That is, if there are $N_{sel}$ original instances that the model answered correctly, the metric is calculated by:
$\text{Conditional}=\frac{1}{\sum_{i=1}^{N_{sel}}k_{i}}\sum_{i=1}^{N_{sel}}\sum_{j=1}^{k_{i}}\texttt{correct}(X_{j}^{s})$ (2)
Hierarchical Consistency
proposed by Wu et al. [66, originally named “Hierarchical Consistence Accuracy”] measures a stricter form of accuracy relative to the previous ones, as the proportion of time the model correctly answers the original unsubstituted instance and all of its corresponding hypernym-substituted instances. Using our notation, this is measured as:
$\text{HC}=\frac{1}{N}\sum_{i=1}^{N}\texttt{correct}(X_{i})\prod_{j=1}^{k_{i}}\texttt{correct}(X_{j}^{s})$ (3)
All of the metrics incorporate robustness to negative samples (using the correct() function). Conditional Accuracy is stricter than Overall, ruling out cases where the model succeeds at higher level categories without correctly answering questions about the target object. HC requires the model to answer all questions about a hypernym chain correctly, being the strictest measure. That is, if the model fails to answer questions about canines then all dog/wolf/fox questions will be penalized. This is the most faithful to the chain in the reference taxonomy but may be considered overly strict.
3.3 Models
We selected seven LM-VLM model pairs, where the LM has been reported to be the base model that the VLM has been trained on top of, following the approach of [23]. The selected pairs are: (1) Llama-3.1-8B vs. MLlama-3.2-11B [12]; (2) their instruct versions; (3) Vicuna vs. Llava-1.5-7B [32]; (4) Mistral-v0.2-Instruct [21] vs. Llava-Next [33]; (5) Qwen2-7B [68] and Molmo-7B-D [8]; (6) Qwen2-7B-Instruct vs. Llava-OneVision [28]; and (7) Qwen2.5-7B-Instruct [69] vs. Qwen2.5-7B-VL-Instruct [4]. See Appendix A for more details. Since TaxonomiGQA consists of Yes/No questions, we sampled from a constrained probability distribution of Yes and No tokens from the models’ output vocabulary, allowing for surface form variation such as casing and space-prefixing.
3.4 Results
The results are shown in Figure 2: points above the diagonal denote model pairs where the VLM outperforms the LM counterpart, and points below denote model pairs where the LM outperforms the VLM. We observe a consistent trend (with a single exception of Vicuna vs. Llava-1.5) where the VLMs outperform their LM counterpart, even though the presentation of the task was purely linguistic. [^1] For the curious: see Appendix D for how the VL models perform on our text-only QA vs. VQA. We rule out the possibility that the VLMs are performing better due to having been trained on GQA directly by running a control experiment where we give only the question (without the scene description) to the VLMs—if the VLMs have already seen the original GQA data, they may have learned the question-label association. The results (Appendix D) show that most VLMs do not perform substantially above chance. Since the text descriptions are newly introduced in TaxonomiGQA, we can safely rule out the hypothesis that VLMs’ improvements are due to having been trained on GQA.
Accepting the trend of VLMs outperforming LMs on TaxonomiGQA as a genuine improvement, we conduct analyses that aim to explain this result more. When we are not analyzing all model pairs, we focus on Qwen 2.5-Instruct vs. Qwen 2.5-VL-Instruct, since the performance gap between VLM and LM was the most salient with this pair especially for the stricter metrics (conditional and HC).
Figure 2: Performance of VLM-LM model pairs on TaxonomiGQA (Section 3) and TAXOMPS (Section 4.1). Points above the line indicate that VLM outperforms LM.
4 H1: VLMs’ taxonomic knowledge aligns better with reference taxonomy
One possible reason that VLMs are performing better on TaxonomiGQA may be due to the difference in their underlying (task-agnostic) taxonomic knowledge, and in particular, in a way that better aligns with the reference taxonomy used to create the hypernym-substituted questions. We test this hypothesis about taxonomic knowledge difference in three different ways: (1) through a QA task that directly elicits taxonomic judgments; (2) through an analysis of the hierarchical organization of concepts in the models’ representation space; and (3) through similarity analysis on the embeddings.
4.1 Directly eliciting taxonomic judgments through Taxonomic Minimal Pairs (TAXOMPS)
Since TaxonomiGQA presupposes taxonomic knowledge rather than eliciting it directly, we first check whether VLMs and LMs differ in their ability to directly answer questions about taxonomic relations. To this end, we introduce TAXOMPS (Taxonomic Minimal Pairs), a dataset which consists of questions of the form “Is it true that a $\bm{C_{1}}$ is a $\bm{C_{2}}$?” where $\bm{C_{1}}$ (cat) and $\bm{C_{2}}$ (feline) are concepts that are in a hypernymy relation, and negative samples where $\bm{C_{2}}$ is replaced by a concept that is not the hypernym of $\bm{C_{1}}$ (vehicle), following Misra et al. [42]. As in TaxonomiGQA, we use 4 negative samples. We construct TAXOMPS directly from the final taxonomy used in our TaxonomiGQA analysis—i.e., 276 total hyponym-hypernym pairs, each coupled with 4 negative samples, yielding 1380 questions. The metric follows Section 3 (specifically, the Overall Accuracy measure, because there is no conditional analog). That is, an instance is considered correct iff. the model answers questions with the hyponym-hypernym pairs (Is it true that a cat is an animal?) with a Yes while answering No to the negative sample questions (Is it true that a cat is a vehicle/fruit/tool/vegetable?).
Figure 2 shows the results. With the exception of Llama-3.1 vs. MLlama 3.2, most VLM-LM pairs perform quite similarly (and well) on TAXOMPS. This suggests that additional VL training does not in general alter the basic taxonomic membership judgments of a language model.
4.2 Lexical representations of taxonomic knowledge
Can the alignment with reference taxonomy be observed representationally, although not by direct elicitation? We test whether the lexical representations in VLMs align better with the reference taxonomy than LMs via their hierarchical organization and hypernym-hyponym embedding similarity.
4.2.1 Hierarchical taxonomic structure
Park et al. [51] propose a method to analyze the latent hierarchical taxonomic structure of an LM, based on ideas including the linear representational hypothesis [38, 52] and causal separability of concepts [65], finding that taxonomic hierarchies (dog < canine < mammal…) are encoded as orthogonalities in LMs’ transformed unembeddings. Therefore, one way we may observe the effect of VL training on the taxonomic knowledge of the LM is via differences in this hierarchical structure.
Method
We applied Park et al.’s method to transform the unembedding space in our models to a space where the inner product between two concepts’ vectors is sensitive to the hierarchical relation between them.Then, we compare VLM-LM pairs in terms of their pairwise cosine similarities between concepts in their unembeddings’ causal inner product space (as established in [52]). In addition, we use the large WordNet hierarchy (a superset of our taxonomy) originally used by [51] to compare the pairwise similarities of concepts in VLM and LM to that of the pairwise path-similarities between concepts in WordNet. These comparisons are done using Representational Similarity Analysis [24], which computes the Spearman’s correlation between two matrices’ (flattened) upper triangular matrices, and take these values as the representational similarity between the three spaces: VLM, LM, WordNet. Greater RSA value between two spaces indicates greater similarity between. To account for potential variance, we sampled 100 subsets (of size 100 $\times$ 100 each) from the larger pairwise matrices and report the mean and standard deviation of the RSA correlations across all subsets.
Results
This analysis (Table 1, left) shows that the hierarchical organization of concepts (as defined by [51]) is mostly shared between the VLM and LM, indicated by the consistently similar RSA scores when comparing VLMs and LMs to WordNet, as well as the high similarity between the VLM and LM when directly compared (all RSA scores $\geq$ 0.95). Interestingly, the Qwen 2.5 and Molmo pairs, the two model pairs that showed the most salient advantage of VLMs in Figure 2 had the lowest VLM-LM RSA scores: 0.95 and 0.96, respectively. However these values are still very high in terms of raw correlation, suggesting that they are still fundamentally similar. The pairwise similarities for VLMs, LMs, and WordNet can be visually inspected in Figure 5 in Appendix E.
Table 1: Left: RSA comparisons of hierarchically sensitive pairwise similarities [51] in the unembedding spaces of VLM-LM pairs, and pairwise path-similarities from the WordNet (wn) Noun Hierarchy. Subscripts show standard deviation (hidden if under 0.01). Right: Differences ($\Delta$) in cosine similarities between positive concept pairs (i.e., in a hypernymy relationship) to those in negative samples from the taxonomy in TaxonomiGQA, computed using VLM and LM static-embedding layers
Minimal Pairs RSA using Park et al. [51] Raw Embeddings
(vlm, wn) (lm, wn) (vlm, lm) $\Delta_{\textsc{vlm}}$ $\Delta_{\textsc{lm}}$ $t$ $p$
Vicuna vs. Llava-1.5 0.43±0.04 0.43±0.04 0.99 0.02 0.02 1.09 0.27
Mistral-v0.2-I vs. Llava-Next 0.42±0.04 0.42±0.03 0.99 0.04 0.04 1.19 0.23
Qwen2.5-I vs. Qwen2.5-VL-I 0.38±0.05 0.39±0.04 0.95 0.03 0.04 -7.51 $<$.001
Llama-3.1 vs. MLlama-3.2 0.40±0.04 0.41±0.04 0.99 0.04 0.04 1.34 0.18
Qwen2-I vs. Llava-OV 0.40±0.04 0.40±0.05 0.99 0.06 0.06 0.82 0.41
Qwen2 vs. Molmo-D 0.38±0.04 0.39±0.04 0.96 0.05 0.05 - -
Llama-3.1-I vs. MLlama-3.2-I 0.40±0.04 0.40±0.04 0.99 0.04 0.04 -0.09 0.92
4.2.2 Embedding similarities of taxonomic relations
Another way in which taxonomic relations can be encoded is via vector similarity—we test the hypothesis that the lexical embeddings (i.e., uncontextualized representations) corresponding to concepts in our reference taxonomy are more similar to embeddings of their hyponyms, compared to embeddings of non-hyponyms in our taxonomy. We compute the similarity between a target concept and four randomly sampled non-hyponym concepts (same as in Section 3.1). Then, we compute the difference between target-hyponym similarity and the average similarity between the target and the negative samples. We test whether this difference is greater in VLMs than LMs, which would mean that VLM embeddings encode hypernym-hyponym relations more similarly than non-hypernym-hyponym relations. This hypothesis is not borne out: Table 1 (right) shows that this holds for no VLM-LM pairs (the significant effect in Qwen2.5-I vs. Qwen2.5-VL-I is in the opposite direction).
5 H2: VLMs are better at deploying taxonomic knowledge
As mentioned in Section 4.1, solving a downstream task presupposes the domain knowledge recruited, and requires this knowledge to be correctly deployed in the context of the specific task. Hence, solving TaxonomiGQA requires (1) taxonomic knowledge and (2) its deployment specifically for scene description-based QA. Our analyses in the previous section did not show convincing evidence in support of the hypothesis that the underlying taxonomic knowledge differs substantially in our VLM-LM pairs. In light of this mostly negative result, we turn to our second hypothesis: VLMs are better at deployment of taxonomic knowledge. To test whether there is a difference when taxonomic knowledge is incorporated into the specific task context, we use contextual similarity of lexical representations and a Principal Component Analysis (PCA) of representations of questions. These analyses let us examine both the contextualized lexical representations as well as the holistic representation of the full question context. We use the Qwen2.5 pair in both analyses.
Data
To control for the confounding effect of the target label (Yes/No) when analyzing contextualized representations, we use a subset of TaxonomiGQA that has the same ground truth label for both positive and negative samples. In our dataset, this only includes cases where the ground truth answer is No. We further filter this dataset to instances where the models got the original, unsubstituted question right, and use the models’ Conditional Accuracy on substituted questions as the target of study. This gives us 37,790 and 40,145 individual samples for Qwen2.5-I and Qwen2.5-VL-I, respectively.
5.1 Contextualized representation similarity
Our first analysis aims to relate the behavioral outcome of a model for each question to the representational structure of the concepts in context. To this end, we investigate the contextualized representations of a target concept in the scene description in terms of their similarity to representations of its hypernym in the question (e.g., There is a doghypo on a yellow surfing board […]. In the scene, are there any mammalshyper?). The quantitative hypothesis is that greater contextualized hypernym-hyponym similarity (e.g., sim(dog, mammal) compared to hyponym similarity with negative samples (e.g., sim(dog, fruit) from There is a doghypo […]. In the scene, are there any fruitsneg?) would predict how well the model can answer the TaxonomiGQA questions. We use the 4 negative samples from TaxonomiGQA, and then fit a logistic regression model to predict model correctness (measured using the correct() function from Section 3.2) using the difference in cosine similarity between hypernym-hyponym, and maximum [^2] We note that taking the average instead of maximum results in substantially weaker trends. cosine similarity of 4 hyponym-negative sample pairs. Here, an odds ratio (of the difference term) being greater than 1 indicates that the similarity of the hyponym to hypernym (relative to negative samples) is more strongly associated with the model correctly answering questions, while the opposite is true if the odds ratio is less than 1. We perform this analysis using representations from every layer in the Qwen2.5 model pair, and take the maximum similarity in cases where the hyponym is mentioned in the scene more than once.
(a) Odds ratio of the difference between cosine similarity of hyponym and hypernym, and max cosine similarity of hyponym and negative samples ($\Delta$) in estimating model correctness, across layers.
[…]
Figure 3: Contextualized representational analysis on Qwen2.5-I and Qwen2.5-VL-I
Results
Figure 3(a) shows the layerwise odds ratios of the difference in similarities between concept pairs (sim-$\Delta$; discussed above), in predicting model correctness, for both models. For most layers, we see odds ratio values that are greater than 1.0, indicating positive association between sim-$\Delta$ and model correctness for both model classes. At the same time, the VLM odds ratios are often greater than those of the LM, with the LM odds ratios sometimes even veering off below the 1.0 level (which would suggest an association of sim-$\Delta$ with wrongness as opposed to correctness. Overall, this suggests that VL training helps establish stronger connections between model representations and behavior in task contexts requiring deployment of taxonomic knowledge.
5.2 Principal Component Analysis (PCA) of question representations
Like the previous analysis, we focus on the distinguishability of hypernym-hyponym relations from non-hypernym-hyponym relations, but consider whether this is captured in the holistic representation of the question context from data used in the previous section. Following Alhamoud et al. [1] (who tested negation sensitivity in VLMs), we take the last hidden state of the final layer of the text decoder to be the summary representation of the full context. Then we ask whether representations of questions that contain a hypernym-hyponym relation (e.g., dog-canine) are separated from representations of questions that contain a non-hypernym-hyponym relation (e.g., dog-bird) via PCA.
Figure 3(b) shows the first two principal components (PCs) of the question representations from the VLM & LM, with hypernym (green) vs. negative sample substitutions (orange) color coded. We see that the two types are largely visually distinct in both models, suggesting that their question representations do encode differences in terms of the taxonomic relations tested. To quantify (linear) separability, we fit a soft-margin support vector machine (SVM) classifier [7] on the first two principal components of the representations extracted from each model separately, and measure its error on the PC-representations—the greater the error, the poorer the separability. We find that the SVM error of the PCs of VLM representations is substantially lower than that of the LM, demonstrating that taxonomic distinctions are more linearly separable in VLM question representations. This complements the results from the previous analysis of contextualized embeddings, and suggests genuine differences in the representational states of the VLM and LM when the task contexts require taxonomic reasoning.
6 Why might vision training help?
Our analyses so far have pinpointed where the meaningful behavioral and representational differences lie in the context of a taxonomic task when comparing a VLM-LM pair. However, we have not discussed why vision training would be beneficial. We present a preliminary investigation here, hypothesizing that visual similarity between members of concepts in a hypernym-hyponym relation is helpful information that VLMs can leverage. Some examples would be the visual similarity of members of equine and horse or root vegetable and radish. Of course, visual similarity will not be informative cues for all such relations, e.g., it would not be very helpful in better understanding the relation between vertebrates and its hyponyms, since there are few salient visual features shared by members of vertebrate (e.g., fish, mammal, amphibian…). This motivates a hypothesis that links visual information to model performance: high visual similarity between members of a hypernym and its hyponym would have a positive effect on model performance on questions probing that relation, but the effect would substantially vary depending on the target concepts.
Method
[…]
Figure 4: Hypernym-specific random effects of image similarity in predicting VLM accuracy on TaxonomiGQA. Greater values indicate closer association of visual similarity to model accuracy. Bar colors indicate percentage of hypernym-hyponym pairs that have above-median similarity.
Results
[…]
7 Conclusion
By building TaxonomiGQA, a text-only QA dataset that requires taxonomic understanding, we identified an interesting performance gap between VLM and their LM counterparts. That is, most VLMs consistently outperformed LMs under all metrics we adopted, despite this task being purely text based. We set out to pinpoint the source of this gain in VLMs. The first set of findings show that both behaviorally and representationally, there was no substantial difference between VLMs and LMs in their taxonomic knowledge, corroborating the general implications of [71, 64] that additional vision training does not fundamentally restructure the underlying knowledge. However, our second set of analyses show: (1) VLMs’ contextual representation similarity of concepts in taxonomic relation in higher layers better predict success on TaxonomiGQA, and (2) VLM representations of questions containing taxonomic relations and questions that do not are better linearly separable, suggesting that VLMs have an advantage over LMs in adequately deploying taxonomic knowledge. We furthermore conducted a preliminary investigation on why vision training helps, testing the hypothesis that visual similarity of members in the extension set of hypernym/hyponyms help VLMs learn more useful representations of these words for taxonomic tasks. The results showed that VLMs’ behavioral success on TaxonomiGQA can be predicted by visual similarity between members of concepts in a taxonomic relation, and the prediction strength is modulated by the visually cohesion of the hypernym.
Limitations and future work
Our statistical analyses do not provide causal evidence for the relation between behavior on TaxonomiGQA and the analyzed representations. Gaining causal evidence would require analyses more closely tied to the training data and objective, which is challenging due to the scale of the models as well as the scarcity of open data models. Nevertheless, we are excited to build on the dataset and methods presented here to conduct deeper analysis addressing the causal question in future work. Furthermore, our SVM-based separability analysis only sheds light on taxonomic distinctions that are linearly encoded, leaving room for future work to extend this to non-linear separability.
Acknowledgments
We thank Yukyung Lee for her advice on creating better visualizations and Mahir Patel for earlier discussions on constructing the dataset, as well as for their general support. Pilot experiments for this work were conducted with the support of the PaliGemma Academic Program GCP Credit Award from Google awarded to the team. An in-person workshop partially dedicated to this work was funded by the Royal Netherlands Academy of Arts and Sciences (KNAW) Early Career Partnership. We acknowledge that the computational work reported on in this paper was performed on the Shared Computing Cluster which is administered by Boston University’s Research Computing Services, as well as on the burrata machine at TTIC.
Appendix A Selected Model Pairs
Table 2 shows a list of model pairs used in this work, along with their metadata – parameters, modality, huggingface identifier, etc.
Table 2: Overview of model configurations used in our experiments, including size, modality, and model training details. The “Training Details” columns indicate whether the model was pretrained on GQA, trained with video data, or instruction-tuned. A checkmark (✓) denotes the presence of the corresponding training signal, (✗) indicates its absence, a bold question mark (?) represents unknown or unclear training status, and a blank cell indicates that the category is not applicable. Hugging Face identifiers are provided for reproducibility.
Model Size Modality Training Details HF Identifier
GQA Pretrained Video Involved Instruction Tuned
Qwen2 7B L Qwen/Qwen2-7B
Molmo-D 7B VL allenai/Molmo-7B-D-0924
Llama-3.1 8B L meta-llama/Llama-3.1-8B
MLlama-3.2 11B VL ? ? meta-llama/Llama-3.2-11B-Vision
Vicuna 7B L lmsys/vicuna-7b-v1.5
LLaVA-1.5 7B VL llava-hf/llava-1.5-7b-hf
Qwen2-I 7B L Qwen/Qwen2-7B-Instruct
LLaVA-OV 7B VL llava-hf/llava-onevision-qwen2-7b-ov-hf
Mistral-v0.2-I 7B L mistralai/Mistral-7B-Instruct-v0.2
LLaVA-Next 7B VL llava-hf/llava-v1.6-mistral-7b-hf
Llama-3.1-I 8B L meta-llama/Llama-3.1-8B-Instruct
MLlama-3.2-I 11B VL ? ? meta-llama/Llama-3.2-11B-Vision-Instruct
Qwen2.5-I 7B L Qwen/Qwen2.5-7B-Instruct
Qwen2.5-VL-I 7B VL Qwen/Qwen2.5-VL-7B-Instruct
Appendix B Extended Related Work
Multimodal Semantic Representations in Humans and Language Models.
A central question in cognitive science and linguistics is how humans integrate perceptual and linguistic signals to form generalizable mappings from semantic and conceptual knowledge to language. Research exploring the cognitive and neural underpinnings of such knowledge supports the idea that language learning and processing is inherently multimodal, grounded in visual, motor, and affective experience [60, 62]. At the neural level, conceptual knowledge is proposed to be coordinated by a transmodal “semantic hub,” allowing humans to flexibly attend to the modality that provides the most informative cue in context and to abstract over modality-specific input [54, 53]. Several NLP tasks now commonly employ multimodal representations [5], notably image captioning [58, 37] and visual commonsense reasoning [72, 30]. In embodied agents, linking physical actions to explicit linguistic representations has been shown to facilitate more effective concept learning [35, 6].
Computational representations can be optimized by identifying and exploiting semantic structure shared across modalities. Models trained on different modalities and objectives may converge on similar representations as they grow in size, forming a “platonic” structure shaped by statistical correlations across input that is modality agnostic [19]. Unified representations and architectures have been argued to better support multimodal processing and reasoning by reshaping how models reference and access perceptual and linguistic features, both reflecting the “semantic hub” structure found in humans and partially mitigating common biases found in unimodal models [59, 10, 67]. These methods can enable the implicit grounding of language in perceptual features such as spatial awareness and sound, even in text-only models [46].
For vision and language modalities, the Visual Question Answering (VQA) task [3] has inspired work on joint language and image understanding using on compositionality, consistency metrics, and knowledge-enriched prompting [22, 11, 9]. Focused benchmarks like VALSE [50], which tests models’ ability to ground linguistic phenomena in the visual modality, and interpretability methods such as MM-SHAP [48] and CC-SHAP [49], measure how VLMs integrate and rely on visual versus textual information. Findings show that VLMs often underuse visual input for reasoning, yet rely on it more heavily for generating explanations, especially in chain-of-thought (CoT) settings. These findings highlight that contributions of each modality in VLMs are uneven and task-dependent, challenging assumptions of uniform multimodal integration. An open question thus remains as to whether multimodal training indeed changes conceptual content, or instead how that content is accessed and applied. Our research explores this in a unique setting where VLM/LM minimal pairs share the textual component.
Appendix C Negative Sampling Details
Table 3: Nine question types in TaxonomiGQA. Each question type is illustrated with an example and the total number of instances of that category. Question types ending in "C" have "no" as the correct answer ("C" stands for counterfactual); all others have "yes" as the correct answer, consistent with the design of GQA.
Question Type Sample Question Counts
exist Are there any dogs? 29030
existAttr Are there any boats that are white? 16405
existAttrNot Are there dogs in this scene that are not white? 15300
existAttrC Do you see dogs that are white? 16010
existAttrNotC Do you see a fork that is not silver? 16440
existThat Are there any tables in the picture that are wooden? 20435
existThatNot Is there a television in the image that is not off? 4120
existThatC Is there a boat that is green? 19985
existThatNotC Is there a watch in the image that is not on? 3670
existMaterial Do you see a fence that is made of wood? 1750
existMaterialNot Is there a bench that is not made of wood? 1465
existMaterialC Are there any lace tablecloths? 1650
existMaterialNotC Are there forks that are not made of metal? 1760
After dataset filtering, we identified 32 types of questions. 19 of them were excluded because substituting the object in the question with one not present in the scene could result in presupposition failures. For the remaining types, we sampled four negative objects for each question based on the following three criteria: the sampled argument is (1) not present in the scene, (2) not in the original argument’s hypernym chain, and (3) associated with the same set of attributes as the original arguments defined in GQA metadata. Due to inconsistencies and errors in the GQA metadata, we manually verified the attribute matches to ensure the naturalness and validity of each substitution. This process resulted in a final dataset consisting of 13 question types and reduced our taxonomy to 126 unique chains. Details of question types, examples, and statistics can be found in Table 3.
Appendix D VQA vs. Text and a Question-only Control
Table 4 shows results from evaluating VLMs on the original GQA questions across different formats: (1) the original VQA setup, conditioned on an image; (2) the Text setup, where they are conditioned on the scene description; and (3) a Question-only control where we condition them only on the question, without any context.
While it is difficult to compare between the VQA and the text setup, we see stark differences in the absolute values of the accuracies. The VLMs seem to answer the (positive sample, unsubstituted) questions with very high accuracy (sometimes near-perfect) relative to their performance on the subset of the VQA task we have used in this work. Next, the VLMs are substantially worse at the question-only baseline than they are in the text version, often times being closer to chance (50%). This question-only control is especially relevant for any potential concern readers might have about VQA data being present in the models’ training set. Since models are largely worse off at these relative to the text version of the dataset, the potential presence of VQA in the model’s training set is of little concern. One interesting observation here is that MLlama-3.2 (non-instruct tuned) performs similarly at the question only task and at the VQA task. This could suggest that it might not have been trained on VQA after all.
Table 4: Accuracies of VLMs on GQA questions when evaluated using standard VQA-based setup (i.e., with images) vs. Text (i.e., with scene descriptions), as well as a Question-only control (No image and no text). Evaluation data consists only the positive sample version of the dataset with unsubstituted questions taken verbatim from GQA [18]. Chance performance is 50%.
Model VQA Text Question-only
Molmo-D 0.79 0.89 0.52
MLlama-3.2-I 0.79 0.92 0.58
MLlama-3.2 0.63 0.91 0.60
Llava-1.5 0.78 0.95 0.53
Qwen2.5-VL-I 0.81 0.98 0.49
Llava-Next 0.84 0.98 0.52
Llava-OV 0.87 0.99 0.60
Appendix E RSA Heatmaps
We depict heatmaps showing the pairwise cosine similarities computed for the transformed Unembedding vectors of the Qwen2.5 model pair, as well as pairwise path-similarity from WordNet, in Figure 5. The LM and VLM matrices look quite similar, while the WordNet matrices are more sparse, showing clearer depiction of hierarhical structure. We computed similar plots for all other models but left them out due to large file sizes. Full plots can be viewed on https://github.com/tinlaboratory/taxonomigqa.
Figure 5: Pairwise similarities between concepts in WordNet, and the transformed unembedding spaces in Qwen2.5-I (LM) vs. Qwen2.5-VL-I (VLM) (computed using Park et al. [51]’s method), across all pairs.
Appendix F More Details about Visual Similarity Analysis
To compute visual cosine similarity between two nodes–a leaf node object (e.g., dog) and one of its hypernyms (e.g., vertebrate)– we first needed a sufficient number of images for both. We used images from THINGS [15], a dataset with 26,107 high-quality, manually curated object-centric images of 1,854 diverse object concepts. Since the taxonomy in THINGS is more coarse-grained than ours, we aligned the taxonomies through the following steps: (1) Identify intermediate nodes that are missing in THINGS (e.g., vertebrate); (2) Collect leaf node objects present in THINGS and prompt a large language model (GPT-4o and Gemini 2.5 Pro) to identify which of them can be classified under the given intermediate node (e.g., vertebrate); 3) Manually verify the correctness of the selected objects. After aligning the taxonomies, we obtained visual representations for each node in our taxonomy from the Qwen 2.5VL-7B Instruct model. To do so, we modified the model’s forward pass to extract hidden states immediately after the merger.ln_q RMSNorm layer within the Qwen2_5_VLPatchMerger module, and before the merger.mlp layer. These intermediate hidden states served as patch-level embeddings, which we mean-pooled to produce a 1280-dimensional representation for each image. We then computed cosine similarities between the visual representation of the leaf node (e.g., dog) and each of its hypernyms (e.g., vertebrate) by taking the mean of all image embeddings for the intermediate category node—similar to the prototype approach in [27]—and compute pairwise cosine similarities with each image from the leaf node.
Appendix G Experimental Resources
Dataset filtering for TaxonomiGQA was performed using multithreaded processing across 8 CPU cores and completed in approximately 3 hours. Negative sampling was carried out on a single CPU core and took approximately 5 minutes.
Model Inference was conducted using vLLM[26]. Vision tasks were processed on a single NVIDIA A40 GPU (48GB) over 3 hours, while text-only tasks were run on two NVIDIA L40 GPUs (48GB each) for approximately 1.5 hours. Image representation extraction for Qwen2.5VL was also performed on a single A40 GPU and completed in roughly 2.5 hours. Static embeddings were computed in under 10 minutes on an L40 GPU.
TAXOMPS, RSA on unembedding layer vectors, contextualized representational similarity analysis, and PCA analysis were conducted on a single NVIDIA RTX6000 Ada (48GB) GPU, and took a total of 1 hour, 1.5 hours, 4 hours, and 1 hour, respectively. Representation extraction and TAXOMPS behavioral analyses were performed using the minicons library [40]. All plots were produced using the ggplot2 library in the R programming language.
We estimate that the total compute cost, including preliminary and unsuccessful experiments, is approximately 3x the sum of the runtimes reported above.
Appendix H License Information
The original GQA dataset was released under CC BY 4.0 and we downloaded the dataset from https://cs.stanford.edu/people/dorarad/gqa/download.html. We follow this and release TaxonomiGQA under the same license, CC BY 4.0.